OpenStudy (anonymous):

Solve for a: A= 1/2h(a+b)

OpenStudy (anonymous):

A=bh/(2-h)

OpenStudy (anonymous):

mulitply a by 2 and distribute h 2A=ha+hb subtract ha 2A-ha=hb extract the a a(2-h)=hb then divide both sides by (2-h) a=hb/(2-h)